OpenStudy (anonymous):

When x = –3 find the value of x3+ 2x2.

hartnn (hartnn):

just plug in x=-3 x^3 +2x^2 = 2^3 +2 (2)^2 whats 2^3 =... ? 2^2=.... ?

OpenStudy (anonymous):

At the market, Fernando weighs his fruit to the nearest 10 grams. He weighs a mango as 260 grams. Complete the statement in the answer space.

hartnn (hartnn):

did u get earlier question ?

OpenStudy (anonymous):

yap

hartnn (hartnn):

so, which statement you need to complete ?